OpenStudy (anonymous):

Please help, I am not good at system of equations at all, I'll give you a medal lol Steve and Mark both received money for their birthdays. Together, they earned a total of $300. Steve earned $50 less than Mark. How much did each boy earn? a. Write a system of two equations with two variables to model this problem: b. Use the substitution or the elimination method to solve the system

OpenStudy (mathmale):

So the guys both rec'd $ for the birthdays? And then they went out and earned a total of $300? ??? Please fix up the language so that it reads more logically.

OpenStudy (mertsj):

Let m = mark's money Let s =steve's money s+m=300 s=m-50

OpenStudy (mertsj):

Replace s in the first equation with m-50 from the second equation and solve.

OpenStudy (anonymous):

@mathmale they received $300 together(as a whole), but Steve earned $50 less than Mark did.

OpenStudy (jdoe0001):

\(\large {\begin{array}{llll} s=Steve\qquad m=Mark\\ \quad \\ \quad \\ \textit{Together, they earned a total of }\$300\implies &s+m=300\\ \quad \\ \textit{Steve earned } \$50 \textit{ less than Mark}\implies &s={\color{red}{ m-50}} \end{array}\\ \quad \\ \bf \implies s+m=300\implies {\color{red}{ m-50}}+m=300}\) solve for "m" to see how much Mark earned to know how much Steve earned, well, he got Mark's less $50, so whatever Mark got, minus 50

OpenStudy (anonymous):

Do you have any idea how I would graph that? I'm completely lost @jdoe0001

OpenStudy (phi):

the question asks in part (b) to solve the problem using either substitution or elimination. why do you want to graph it ? If you really want to graph it, I would rename your variables x and y. (It does not matter which is which, just make a note of what x and y stand for) For example, if mark is x and steve is y, rewrite s+m=300 --> y + x =300 s=m-50 ----> y= x - 50 you can rewrite the first equation y+x=300 in slope intercept form: y= -x + 300 the 2nd is already in slope intercept form: y= x-50 now you need to know how to plot lines.